SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Operation Meeting (04/08/08)

Attendees: Rahmat Rahmat, Kim Hojeong.

Meeting Information

 Date      :  April 8th, 2008
 Time      :  2:00pm - 3:30pm PST
 SLAC Room :  Madrone conf room, ROB 
 
 Dial-in   :  1-510-665-5437, pass code 6649#

--- Minutes ---

Operations

  • No problem. We stopped data taking on Monday ~12:45pm.

EMT

DCT

GLT

Online

L3

DQM

  • L3T Boost plot during Sat Apr 5th Swing shift. Rahmat will find out what are the definitions of these variables.
  • TSF occupancy in Cells in SL1 has small peak near 50. It showed up around Jan 21/22. We think due to high background. On Jan 21/22, we had power glitch and very high background. Why it lasts when we have low background is a puzzle.
  • Rahmat checked stripcharts and everything is OK.

Trickle injection monitoring

Trigger Configuration

AOB

  • Next week, the presumably last meeting, we will have cookies and drinks.

This page was created by Kim Hojeong; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.

Last Modified: April 8th, 2008.